### v1

# VecTrans: LLM Transformation Framework for Better Auto-vectorization on High-performance CPU

Abstract.
Large language models (LLMs) have demonstrated great capabilities in code generation, yet their effective application in compiler optimizations remains an open challenge due to issues such as hallucinations and a lack of domain-specific reasoning. Vectorization, a crucial optimization for enhancing code performance, often fails because of the compiler’s inability to recognize complex code patterns, which commonly require extensive empirical expertise. LLMs, with their ability to capture intricate patterns, thus providing a promising solution to this challenge. This paper presents VecTrans, a novel framework that leverages LLMs to enhance compiler-based code vectorization. VecTrans first employs compiler analysis to identify potentially vectorizable code regions. It then utilizes an LLM to refactor these regions into patterns that are more amenable to the compiler’s auto-vectorization. To ensure semantic correctness, VecTrans further integrates a hybrid validation mechanism at the intermediate representation (IR) level. With the above efforts, VecTrans combines the adaptability of LLMs with the precision of compiler vectorization, thereby effectively opening up the vectorization opportunities. Experimental results show that among all 50 TSVC functions unvectorizable by Clang, GCC, and BiShengCompiler, VecTrans successfully vectorizes 23 cases (46%) and achieves an average speedup of 2.02x, greatly surpassing state-of-the-art performance.
1. Introduction
Vectorization is an critical compiler optimization that transforms scalar operations into vectorized forms to exploit the parallel processing capabilities of SIMD (Single Instruction, Multiple Data) architectures. Vectorization dramatically improves computational throughput by simultaneously processing multiple data elements as a single vector operation, making it indispensable for performance-critical domains such as scientific computing, machine learning and multimedia processing. The theoretical foundation of vectorization dates back to Flynn’s taxonomy of parallel architectures(flynn2005very, ), but its practical implementation has evolved alongside advancements in processor design, particularly with the proliferation of SIMD instruction sets such as Intel AVX(IntelAVX2, ), ARM SVE(stephens2017arm, ), and RISC-V V Extension(cui2023risc, ). Despite decades of researching and engineering investments, automatic vectorization still faces serious challenges, including complex data dependencies and diverse code patterns.
Currently, several strategies are available to address non-vectorized loops, including manually modifying the source code, using intrinsic instructions, and applying targeted vectorization optimizations. While these methods can be effective, they each come with their own set of drawbacks. Manual source code modifications (e.g., rewriting loops into SIMD-friendly structures) preserve readability through standard C/C++ implementations, enabling cross-platform compatibility via compiler portability. However, such hand-tuning demands significant developer effort and risks introducing subtle errors like incorrect dependency handling. Compiler intrinsics, e.g., Intel’s _mm512_load_ps, provide precise hardware control, but sacrifice portability by binding code to specific instruction sets (e.g., AVX-512 intrinsics fail on ARM platforms) and obfuscating algorithmic intent. Targeted optimizations automate some vectorization tasks but struggle with non-affine loops or dynamic control flow.
Even compiler-driven approaches confront fundamental tensions: while vanilla C/C++ code retains platform independence, reliance on compiler auto-vectorization leads to inconsistent results, where LLVM(lattner2004llvm, )’s Loop Vectorizer succeeds in some cases, and GCC’s Vectorizer performs better in others. The core challenge lies in conflicting priorities: maintaining portable source code requires abstracting hardware details, whereas maximizing SIMD utilization demands deep architectural awareness (e.g., register widths, masking support). This dichotomy forces developers into a lose-lose choice: either accept suboptimal performance with portable scalar code or lock into vendor-specific intrinsics for marginal gains. Moreover, the inherent limitations of heuristic-driven nature of modern compiler vectorization often fails to fully exploit the available SIMD parallelism, leaving significant performance potential untapped.
These limitations collectively highlight the need for a paradigm that transcends the manual-vs-automatic divide—one that preserves code portability while achieving optimizations close to hand-tuned intrinsics. Our work addresses this challenge by leveraging LLMs to infer architectural constraints from code semantics, enabling automated vectorization that adapts to both program logic and hardware capabilities without compromising readability or portability. Pretrained on vast code corpora, LLMs exhibit emergent capabilities in understanding complex program semantics and generating context-aware transformations. Unlike rule-based compilers, LLMs can reason out implicit parallelism opportunities in loops with irregular structures, such as those containing conditional branches or indirect memory accesses. Our approach employs a multi-feedback architecture that simultaneously analyzes data dependencies, predicts vectorization viability, and generates optimized loop variants that can maximize the potential of LLM-driven code transformation.
However, naively applying LLMs to code optimization presents new challenges. Firstly, LLM is a category of probabilistic models proficient in solving some combinatorial explosion problems. Nevertheless, it suffers from serious hallucinations especially in the tasks related to strong reasoning like mathematical or code generation problems. These unresolved questions make its capability unreliable(zhou2024larger, ; mirzadeh2024gsm, ) if utilized in our studied auto-vectorization optimization. To make LLM’s ability more robust and deterministic, a solid framework usually called AI agent is essential to enhance the interaction of LLM with other robust and precise tools like traditional compilers. Secondly, how to generate both syntactically and semantically correct code is usually a fundamental requirement for auto-vectorization in traditional compilers, but the situation is not suited to naive LLMs. To overcome these challenges, we designed a verification-aware multi-feedback iterative refinement (VAMF-IR) framework, also able to viewed as an AI Agent, for auto-vectorization optimization. The iterative refinement method is proposed by Wilkinson(wilkinson1965rounding, ) in linear system solving problems. Simultaneously, it is firstly introduced in LLM by Madaan(madaan2023self, ) to remarkably enhance the capability of naive LLMs using self-refine mechanism Correspondingly, VAMF-IR is divided into three phases: 1) generate the candidate transformed source code using LLM model; 2) verify semantic equivalence between the original and transformed source codes utilizing unit test and, successively, formal verification with a symbolic execution engine; 3) invalid candidates trigger model retraining with counterexamples, progressively improving vectorization accuracy. This feedback loop enables the system to surpass the coverage limitations of static analysis while maintaining correctness guarantees, marking a critical advancement over existing compiler-based approaches.
In summary, the paper makes the following contributions:
  • LLM-guided source code transformations for vectorization: We introduce a novel approach that leverages LLMs to enhance vectorization opportunities through source code transformations. Our proposed method ensures cross-platform versatility while generating more human-readable code.
  • LLM-compiler collaboration paradigm: We explore a new paradigm that combines the inference capabilities of LLMs with the optimization strengths of compilers. This approach enables more accurate and higher-performing vectorization.
  • Cross-platform verification: Our method ensures correctness across platforms by performing formal verification at the intermediate representation (IR) level, eliminating the need for platform-specific validation.
  • Effective results: The evaluations show that VecTrans can effectively extend the vectorization cases on the representative TSVC_2, and further achieves significant performance improvements.
void s1113(int iterations, int LEN_1D, float* a, float* b)
for (int nl = 0; nl < 2*iterations; nl++) {
for (int i = 0; i < LEN_1D; i++) {
a[i] = a[LEN_1D/2] + b[i];
(a) Original code
void s1113_opt(int iterations, int LEN_1D, float* a, float* b) {
for (int nl = 0; nl < 2*iterations; nl++) {
int mid = LEN_1D / 2;
float temp = a[mid];
for (int i = 0; i < mid; i++) {
a[i] = temp + b[i];
a[mid] = temp + b[mid];
temp = a[mid];
for (int i = mid + 1; i < LEN_1D; i++) {
a[i] = temp + b[i];
(b) Optimized code
Figure 1. An example from TSVC_2. LLVM fails to vectorize the original code in (a) due to dependence analysis, restructuring the code in (b) allows the compiler to successfully perform auto-vectorization.
2. Background and Motivation
2.1. Background
Modern compilers generally employ two primary vectorization strategies, namely loop vectorization and Superword Level Parallelism (SLP) vectorization(larsen2000exploiting, ). Loop vectorization, the more established approach, transforms entire loop iterations into vector operations by exploiting cross-iteration parallelism. However, this method requires an extra verification given that loop-carried dependencies are absent. In contrast, SLP vectorization operates at the basic block level, identifying and grouping isomorphic scalar operations within loop bodies or straight-line code into vector instructions (porpodas18lslp, ; porpodas18vwslp, ; porpodas19snslp, ). Both vectorization approaches face significant limitations when dealing with real-world complex codes. Loop vectorization inherently fails to handle loop-carried dependencies situations where iteration i+1 depends on results from iteration i (e.g., recurrence relations like prefix sums). SLP vectorization, while less constrained by cross-iteration dependencies, struggles with non-isomorphic operation patterns.
Several studies have explored the application of LLMs in code optimization. LLM-Vectorizer(10.1145/3696443.3708929, ) demonstrates notable performance by transforming source code into vectorized implementations through intrinsic instructions. However, it exhibits three limitations: 1) inadequate handling of loop-carried dependencies that hinder effective parallelization, 2) generation of platform-specific constructs that compromise portability across architectures, and 3) production of low-readability code lacking maintainability. In parallel, Self-Refine(madaan2023self, ) employs iterative LLM-based refinement to progressively enhance code quality. Nevertheless, this approach suffers from insufficient integration with compiler semantics, resulting in suboptimal optimizations due to the absence of compiler-guided metrics and target-specific optimization objectives. Both paradigms reveal fundamental challenges in integrating LLMs with compiler optimization frameworks.
Figure 2. Overview of VecTrans. When a user submits source code requesting vectorization, if the target function is identified as non-vectorizable, the framework initiates an iterative optimization process. Each iteration comprises two phases: feedback collection and refinement. During the feedback phase, the framework aggregates three categories of diagnostic information: self-evaluation feedback from large language models, static analysis feedback from compilers, and runtime validation feedback. The refinement phase then leverages these diagnostic insights to perform semantics-preserving transformations aimed at enhancing vectorization potential. The framework evaluates the optimized function’s vectorizability, proceeding to subsequent iterations if residual vectorization barriers persist with remaining optimization potential. This iteration process terminates either upon successful vectorization achievement or upon reaching optimization convergence.
2.2. Motivation
While compilers have made significant progress in auto-vectorization, they still face fundamental limitations in recognizing and transforming certain code patterns. Even for recently pioneering work related to LLM-Vectorizer(10.1145/3696443.3708929, ) which tries to integrate LLM, compiler and verification into an AI agent framework, it also loses many optimization opportunities possibly due to the non-optimal interaction manner among different agent components.
To better illuminate the motivation of our work, a representative example is illustrated in Figure 1. The original source code as given in Figure 1(a) cannot be vectorized with traditional compilers even on O3 mode for two reasons: (1) when i=LEN_1D/2, a[LEN_1D/2] is updated, which results in a read-after-write (RAW) dependency in subsequent iterations, and (2) simultaneous reading and writing an array a within the loop makes vectorization unsafe, causing the compiler to reject the transformation. Upon further analysis, it is found that a[LEN_1D/2] is a variable and is updated only once, making it loop-invariant when the loop is split into two separate sections. Unfortunately, LLM-vectorizer(10.1145/3696443.3708929, ) also fails to generate the code able to be vectorized even after several iterations. This is mostly because the generated SIMD intrinsic code cannot pass the verification and also cannot be corrected. It indicates that we need to seek for more opportunities on source code level rather than low-level code like intrinsic or assembly, because general LLM are usually well trained on source code datasets.
Motivated by the above experimental and analyzed results, we believe that the interaction manner among LLM, compilers and other tools makes a difference to overall optimization results for auto-vectorization. Corresponding, in our current work, VecTrans framework is developed to resolve or massively improve the interaction behaviors of different components in source code levels. It integrates different advantages for different components including LLM, compilers and verification tools into a multi-feedback framework - VAMF-IR. It means that LLM is responsible for finding creative opportunities because of its high-level correlation analysis strength, while compilers and verification tools are usually conservatively designed, they are more appropriate to provide some accurate feedback information. Besides, they are more suitable to compile source code precisely on backend SIMD instructions than LLM. The further description about VecTrans are presented in the next section.
3. Design
VecTrans is an iterative refinement LLM optimization framework that uses feedback-refined iteration structures(madaan2023self, ). The feedback-refined iteration is a cyclical optimization process comprising two phases: the feedback phase and the refine phase. During the feedback phase, information relevant to the optimized function is gathered, referred to as feedback information. This feedback is utilized to assess the correctness of the optimized function and to identify the potential for vectorization. In the refine phase, this feedback is applied to update and enhance the optimized function. The feedback information is derived from three sources, as illustrated in Figure 2:
  • Source 1: Self-Feedback. The basic feedback information comes from the capabilities of the LLM itself.
  • Source 2: Compiler Feedback. Feedback information from the compiler, including compilation information and vectorization information. This information is used to help the LLM determine whether the currently generated optimized code can be compiled and vectorized successfully.
  • Source 3: Verification Feedback. The feedback information is used to check whether the generated optimized code can pass the unit tests and formal verification. This helps the large model determine the semantic correctness of the optimized code.
3.1. Source 1: Self-Feedback
The self-feedback information is derived from the analytical capabilities of the LLM. Specifically, the model is leveraged to assess various aspects of the currently optimized code, including lexical correctness, syntactic accuracy, semantic integrity, and the effectiveness of vectorization. By conducting a comprehensive evaluation, the LLM generates valuable insights that inform the next steps in the optimization process. The analysis results produced by the model subsequently guide the refinement phase, where the code undergoes further enhancement based on the feedback provided.
Figure 3. The compiler provides feedback through lexical/syntactic errors, semantic unit test results, and vectorization insights via Rpass flags. This multi-source diagnostic evaluates code validity and vectorization potential.
3.2. Source 2: Compiler Feedback
Compiler feedback information consists of three parts as shown in Figure 3. In the first part, it conducts a compilation-error check on the optimized code, capitalizing on the error-related information provided by the compiler. In the second part, it performs a semantic-equivalence analysis by comparing the source code and the optimized code through unit test. In the third part, the Rpass and Rpass-analysis parameters of the compiler output vectorized information about the optimized code, including the positions of the loops that were successfully vectorized, the positions of the loops that were not successfully vectorized, and the reasons why the vectorization was not successful. These three parts of information can help the LLM determine whether the current code is optimized and fully vectorized while ensuring semantic correctness.
Figure 4. Verification feedback ensures code correctness through dual mechanisms. Unit testing verifies semantic equivalence by comparing execution outputs, though constrained by input coverage limitations. Formal verification complements this approach by mathematically proving behavioral equivalence through specification-based analysis, establishing multi-layered correctness guarantees.
3.3. Source 3: Verification Feedback
The verification feedback is divided into unit test information and formal verification information shown in Figure 4:
1. Unit test: This tests the syntax and semantic correctness of the code. The generated optimized code must first pass compilation and ensure that its semantics are consistent with the source code. In this phase, the LLM is used to generate unit test code, which includes generating simple inputs and validating outputs. Unit tests allow the correctness of the optimized code to be verified within the scope of the covered inputs. To ensure efficiency, unit test code is generated only once at the beginning and reused in subsequent iterations. To ensure the correctness of the unit test code, it is validated by replacing the optimized code. First, the original function replaces the optimized function, and the output of the optimized function must match the original function, as the two functions are identical at this point. If the return value is ’void’, an empty function is used to replace the optimized code. If the return value is not ’void’, the return values of the original code are bitwise negated. At this point, the outputs of the original and optimized functions must differ; otherwise, the unit test code is considered incorrect. There is a flaw in this approach, as if the original function is empty, a false positive might occur, but this situation is rare, so it is ignored here.
2. Formal verification: After the optimized code passes the unit tests, it must undergo formal verification. Alive2(10.1145/3453483.3454030, ) is used as the formal verification tool. Both the original and optimized functions are compiled into IR and then verified using Alive2. If the formal verification is successful, the code moves to performance testing. If it fails, the process loops back with the addition of error information.
3.4. Combination of Feedbacks
In the refine phase, a comprehensive set of inputs, namely the source code, the current state of the optimized code, the feedback information, and relevant prior knowledge, is integrated. The prior knowledge encompasses concepts such as loop splitting, loop reordering, instruction reordering, conditions for using temporary variables, along with simple illustrative examples like Instruction Splitting (which requires independence between instructions) and Iteration Range Splitting (where the loop should contain phase-dependent patterns).
This phase capitalizes on the feedback from the previous feedback phase. The LLM receives as inputs the feedback data, the source code, and the optimized code. Leveraging this information and the prior knowledge, the model makes informed decisions to update the optimized code. The output of this process is the refined optimized code, which is carefully crafted to maximize performance and adhere to programming standards.
A key aspect of this phase is the constraint of generating only the optimized code, which is encapsulated with specific identifiers. This not only streamlines the output but also enables easier post-processing and integration. By decoupling the feedback and iteration phases, we enhance the LLM’s performance. It allows for more focused processing within each phase, facilitating better organization of the overall optimization process. This, in turn, makes the entire procedure more conducive to automation, as the clear separation of tasks and well-defined input-output relationships simplify the implementation of automated workflows.
3.5. Optimized Case
The framework has the ability to handle complex functions from real-world applications(Skia), as shown in Figure 5. The temporary array is used to temporarily store intermediate results, and loop splitting is used to separate the parts with loop dependencies, so that the remaining loops without dependencies can be vectorized successfully.
void set_points(...) {
for (int i = 0; i < divCount; i++) {
src[i + 1] = divs[i];
int srcDelta = src[i + 1] - src[i];
float dstDelta;
if (srcFixed <= dstLen) {
dstDelta = isScalable ? scale * srcDelta : srcDelta;
} else {
dstDelta = isScalable ? 0.0f : scale * srcDelta;
dst[i + 1] = dst[i] + dstDelta;
isScalable = !isScalable;
(a) Original code
void set_points_opt(...) {
for (int i = 0; i < divCount; i++) {
src[i + 1] = divs[i];
srcDelta[i] = src[i + 1] - src[i];
if (srcFixed <= dstLen) {
dstDelta[i] = isScalable ? scale * srcDelta[i] : srcDelta[i];
} else {
dstDelta[i] = isScalable ? 0.0f : scale * srcDelta[i];
isScalableArray[i] = isScalable;
isScalable = !isScalable;
for (int i = 0; i < divCount; i++) {
dst[i + 1] = dst[i] + dstDelta[i];
(b) Optimized code
Figure 5. An example from Skia
4. Evaluation
4.1. Experimental Methodology
System Configuration. We perform most experiments on a Linux-based platform equipped with a 128-core Kunpeng-920 AArch64 CPU(8MB L1d cache, 8MB L1i cache, 64MB L2 cache, 128 MB L3 cache). For validation, we employ BiSheng Compiler 3.2.0.1(bisheng, ) with optimization flags -O3 -ffast-math, -Rpass=loop-vectorize, -Rpass-analysis=loop-vectorize to enable vectorization diagnostics and feedback.
Benchmarks. We evaluate a wide range of functions from TSVC_2(maleki2011evaluation, ) and real application function from Skia(google2014skia, ). The TSVC_2(Test Suite for Vectorization Compilers) benchmark suite systematically evaluates compiler capabilities in automatic vectorization through 149 test functions representing typical computational patterns. Experimental results reveal that 50 functions (approximately 33.6%) resist effective vectorization by BiSheng Compiler(bisheng, ) and mainstream compilers, primarily due to complex data dependencies, irregular memory access patterns, or unstructured control flows. The classification of these functions is shown in Table 1. These selected cases serve as critical evaluation targets, exposing limitations in current vectorization techniques while providing a standardized framework for analyzing compiler optimization boundaries and guiding algorithmic improvements.
Metrics. We consider the effectiveness of the framework from three dimensions: coverage, acceleration ratio, and iteration rounds. All of the selected functions cannot be automatically vectorized by the compiler. Through the transformation of the framework, some functions can be automatically vectorized. The proportion of these functions, that is, coverage, is our primary consideration. The speedup of the test functions that can be successfully vectorized is another considered metric reflecting the benefits of the framework. Iteration rounds reflect the resources consumed by conducting this framework for optimization, including the number of iteration rounds needed for functions that fail to be automatically vectorized, and reflect the cost of deploying the framework in the compiler.
Parameter Setup. The LLM selected for the experiment is Deepseek-V3(liu2024deepseek, ). The maximum number of iteration rounds is set to 20. When the unit test generates and analyzes whether the function is successfully vectorized (feedback), temperature is set to 1 and temperature is set to 0 in other phases. A higher temperature corresponds to increased output randomness, with the minimum temperature is zero. Set the maximum number of tokens to 4096 and retain the default values for other parameters.
Table 1. The classification of functions that cannot be vectorized in TSVC_2.
Category Function name
Unsafe dependent memory operations (13) s112, s1113, s114, s115, s116, s141, s212, s241, s242, s244, s1244, s281, s293
Could not identify reduction variable (33) s116, s123, s126, s141, s161, s211, s212, s1213, s221, s222, s231, s232, s233, s2233, s235, s2251, s256, s258, s261, s275, s277, s291, s292, s2111, s318, s3110, s3112, s321, s322, s323, s341, s342, s343
Cannot identify array bounds (2) s123, s1161
Could not determine number of loop iterations (5) s161, s277, s332, s481, s482
Instruction cannot be vectorized (2) s31111, s451
Loop contains a switch statement (1) s442
Figure 6. VecTrans generates optimized code using BiSheng Compiler(bisheng, ) on the Neon instruction set with an average speedup of 2.02x, compared with an average speedup of 1.94x for LLM-Vectorizer. The bar marked with a slash indicates that vectorization fails. It can be observed that LLM-Vectorizer and VecTrans can cover different test cases because LLM-Vectorizer is better at processing functions with conditional branches. VecTrans is better at dealing with functions with dependencies between loops.
4.2. Coverage
Our model can successfully vectorize 23 of the 50 TSCV functions, as well as a complex function from skia, coverage reached 46.2%. The coverage of LLM-Vectorizer on Neon instruction sets is only 28.8%. The functions that can be successfully vectorized are shown in Figure 6. According to the different categories shown in Table 1, among which the functions that cannot be vectorized due to unsafe dependent memory operations are most easily transformed by this framework so that they can be vectorized automatically, 10 out of 13 functions can be vectorized successfully. The functions that cannot be vectorized because they could not identify the reduction variable are more difficult to transform. Only 13 of the 33 functions can be vectorized successfully. The remaining four categories are difficult to handle. Only one test case s1161 can be vectorized successfully. Other test cases cannot generate optimization functions that can be vectorized automatically.
Table 2 shows the manner in which the LLM transformation function enables vectorization. Through these transformations, the compiler can more easily identify whether there is a dependency between loops, so that the compiler can vectorize a function that could not be vectorized originally. One of the most effective is the loop splitting technology, which can effectively eliminate many loop dependencies.
Table 2. Code Optimization Techniques and Descriptions
Optimization Description
Loop Splitting Decompose a single loop into multiple independent sub-loops to enable parallel execution. Commonly used for reducing loop-carried dependencies.
Instruction Reordering/Loop Reordering Reorganize instruction execution sequence to eliminate false dependencies, or rearrange loop nesting order to enhance data locality.
Temporal Variable Introduce short-lived variables to store recurrent computation results, reducing dependency caused by array access..
Branch Elimination Remove conditional statements like switch or if.
In addition to the test cases that can be successfully vectorized, there are 27 test cases that cannot be successfully transformed by the framework. The conversion fails due to the following reasons:
  • The number of iteration rounds is exceeded. The upper limit of the number of iterations is set to 20. When the number of iterations exceeds 20, if the LLM still considers that the function at this time cannot be automatically vectorized, the iteration is terminated and the source code is output.
  • The LLM considers that the code has been optimized. In some cases, the LLM will ignore the vectorization information provided by the compiler. As a result, the LLM mistakenly considers that the current function can be automatically vectorized and chooses to end the iteration and output the source code.
  • The LLM considers that there is no optimization space. The LLM considers that the existing loop dependency makes the function unable to complete automatic vectorization through the compiler, and chooses to end the iteration and output the source code.
  • Generate wrong code. Among the 51 test functions, there is only one function s481, and the optimization function generated by the function s481 can pass the unit test and formal verification. However, the semantics of the optimization function is different from the original semantics, mainly because the input coverage of the unit test is limited. In addition, the formal verification tool is incomplete. This situation will output an error code, which is unacceptable, and future work will attempt to deal with it.
The functions that cannot be successfully converted by the framework have the following features: two-dimensional arrays, conditional branches, and function calls. In addition, if the loop dependency cannot be resolved, the loop cannot be vectorized successfully. The main factor is the existence of conditional branches. There are 11 test cases that cannot be transformed by the framework because of the existence of conditional branches. Second, there are 8 functions that cannot be resolved because of loop dependencies.And there are six functions that cannot be vectorized due to the difficulty of memory access analysis caused by arrays. The remaining two test cases are due to function calls, function inlining is required before the next step is carried out, which is difficult to implement at the source code level.
4.3. Speedup and Iteration Rounds
Under the Neon instruction set, a total of 24 functions are successfully vectorized using the BiSheng Compiler(bisheng, ). In comparison to the O3 optimization level, which achieves an average speedup of 2.02x and a maximum speedup of 5.21x (for function s293), two functions exhibit negative optimization outcomes and four functions show minimal or no performance improvement. This degradation in performance can be attributed to the additional overhead introduced by the transformation process, such as the creation of temporary arrays, cache misses, and other inefficiencies, which ultimately outweigh the benefits of vectorization.
The optimization framework typically necessitates a mean of 8.76 iterations per function. Functions exceeding this threshold undergo optimization cycles up to a maximum of 20 iterations, with 9 functions reaching this upper bound. Successful vectorization requires significantly fewer iterations, averaging 4.96. Each iteration round comprises two distinct Large Language Model (LLM) interactions: the feedback phase, where performance data is collected and analyzed, and the refinement phase, where the optimization function is updated based on the feedback.
Table 3. The coverage and rounds under different configurations. The Base Model means that the framework without formal verification, compiler feedback and unit test.
Configuration Coverage Iteration rounds
DeepSeek-V3/VecTrans $\mathbf{46.2\%}$ 8.760
DeepSeek-V3/LLM-Vectorizer(10.1145/3696443.3708929, ) 28.8% 13.39
DeepSeek-V3/Base Model 17.3% 5.412
Qwen2.5-32B/VecTrans 34.6% 13.940
Qwen2.5-72B/VecTrans 38.5% 12.260
DeepSeek-V3/Without Formal Verification 32.7% 8.489
DeepSeek-V3/Without Compiler Feedback 21.2% 5.213
DeepSeek-V3/Without Unit Test 25.0% 9.656
Figure 7. The graph shows the speedup of functions that can be successfully vectorized when the framework uses different LLMs. Columns with slashes indicate that vectorization cannot be successful. Three LLMs, Qwen2.5-Coder-32B-Instruct(hui2024qwen2, ), Qwen2.5-72B-Instruct(yang2024qwen2, ) and Deepseek-V3, are used here. The speedup is of the final generated optimized function compared to the original function using BishengCompiler at -O3 -ffast-math flag.
4.4. Breakdown Analysis
To understand the impact of the different components of the framework on VecTrans, we evaluated the impact of formal verification, compiler feedback, and unit test on the overall framework.
Table 3 shows the coverage and number of iteration rounds under different configurations. It can be seen that when a component is removed, the coverage decreases, and when all three components are removed (Base Model), the coverage reaches the lowest value. This reflects the positive effect of all three components in helping the framework complete code conversion. Compiler feedback has the greatest impact, and formal validation has the least impact.
Compiler feedback helps the entire framework in three ways as follows.
  • Helps LLM determine whether a compilation error occurs. If a compilation error occurs, the optimized function needs to be regenerated.
  • Helps LLM determine whether vectorization is successful. If vectorization has not been performed, LLM needs analyze whether vectorization opportunities exist and provide conversion suggestions.
  • Helps with LLM analysis which transformation can facilitate automatic vectorization by the compiler. The compiler gives the loop position that cannot be vectorized and the reason why it cannot be vectorized.
Therefore, when the compiler feedback is removed, the LLM needs to rely on its analysis capability to determine whether the current optimized function can be automatically vectorized. When the LLM mistakenly considers that the current function can be automatically vectorized, the LLM chooses to end the iteration and output code. In fact, the current function is not successfully vectorized. In experiments, this is common, so when the compiler feedback is removed, the coverage decreases, and the number of iteration rounds decreases, because the LLM considers that the current optimized function has been fully optimized in the first few iteration rounds and selects to output the code.
Unit test and formal validation are similar in that they are used to detect the semantic correctness of the generated code, where unit test is more important than formal verification in that it can intercept more incorrect code because formal verification will fail in some complex situations. The unit test is more applicable, but it cannot cover all cases due to the limited input range. Therefore, formal verification is needed as a supplement.
The removal of different components has an impact on the coverage rate and iteration rounds, but has little impact on the acceleration ratio because the code optimization methods used by the same LLM are similar and the generated code is basically the same. Except that a few functions may reduce the acceleration ratio due to redundant operations, the acceleration ratio remains the same in other cases.
Figure 8. Average speedup of optimized function generated by VecTrans under different configurations
4.5. Sensitivity Studies
We verify the robustness of the framework from three dimensions: model, compiler, and architecture. Only one parameter is changed at a time, and the rest remain unchanged. We have verified that the framework works across different LLMs, compilers, and architectures.
Sensitivity of different models. We evaluate the usability of the framework on three LLMs, Qwen2.5-Coder-32B-Instruct, Qwen2.5-72B-Instruct, and Deepseek-V3. In addition, small-scale tests are performed on smaller models. It is found that the instruction compliance capability of smaller models is poor and the normal iteration process cannot be completed. The Deepseek-R1 model is also tested on a small scale, which can produce correct results in fewer iterations, but takes too long.
The coverage rate and iteration rounds of the three different base models are shown in Table 3. It can be seen that with the decrease of model parameters, the coverage rate decreases and the number of iteration rounds increases. This is because the decrease of model parameters leads to the weakening of model capabilities, so some test functions are difficult to successfully transform. At the same time, more iteration rounds are required, and some test functions cannot output correct results within 20 iteration rounds.
Although the capabilities of the framework are affected by the base model, the coverage of small models is higher than that of large models without framework components. Therefore, the framework is robust on different models and can be migrated to different LLM.
Figure 7 shows the speedup of different model generation functions. The speedup of Qwen2.5-Coder-32B-Instruct is 2.9 times, and that of Qwen2.5-72B-Instinct is 2.61 times. The speedup of Deepseek-V3 is 2.54 times. Even for different LLMs, the generated functions are similar and the speedup is similar. Therefore, this framework may help identify potential compilation optimization points, that is, summarize the characteristics of the generated functions. The optimization direction of automatically vectorized can be found.
Sensitivity of different compilers. Here we explore the impact of different compilers on the optimized functions, and we test the speedup of GCC 14.2.0, Clang 20.0.0, and BiSheng Compiler 3.2.0.1(bisheng, ), respectively. The GCC and Clang versions used here are the latest versions, which represent the best compilers available today. BiSheng Compiler is a compiler optimized for Kunpeng series chips. On Kunpeng 920B, the Neon vector instruction set is used, and the speedup of functions compiled using the flag -O3 -ffast-math on different compilers is shown in Figure 8.
The average speedup of the optimized function is 1.95 times on Clang, 2.07 times on GCC, and 2.54 times on BiSheng Compiler(bisheng, ). The code generated for the s112 function is significantly worse on GCC and Clang than it is on BiSheng Compiler(bisheng, ) because the latest versions of GCC and Clang can vectorize the original s112 function. In s1161 and s242, the speedup of GCC is lower than that of the other two compilers. In s293, the speedup of Clang is lower than that of the other two compilers. This indicates that different compilers optimize functions differently. However, in most cases, the optimization of these functions is effective for all three compilers, and can achieve good results.
Sensitivity of different architecture. Different vector instruction sets also have a great impact on optimized function performance. In the x86 architecture, AVX instruction sets are usually used. In the ARM architecture, NEON instruction sets and SVE instruction sets are used. We will explore the effects of different vector instruction sets on the performance of generating optimized functions on different architectures.
As shown in Figure 8, the average speedup is 2.80 times on the AVX instruction set, 2.02 times on the Neon instruction set, 2.54 times on the SVE instruction set, and on the AVX instruction set, the speedup of the s1161 function can reach 16.8 times.
It can be observed that, on different vector instruction sets, there is no obvious regularity in the speedup of the function, but the effect is better on the AVX instruction set and the SVE instruction set. The effect comes from the length of the vectorized register on the one hand, and the vectorized implementation mode on the other hand.
4.6. Summary of Results
In summary, the framework effectively transforms functions so that functions can be vectorized automatically by the compiler. In our experiments, this framework successfully vectorized 24 of the 51 functions, achieving an average speedup of 2.54 times. At the same time, this framework can improve the code generation capability of LLMs, even small models can guarantee the function generation effect. Through unit test and formal verification, the correctness of the generated code is ensured, so that the framework can be deployed in the compiler. Finally, we verify the effectiveness of the generated optimized code on different compilers and vector instruction sets.
5. Related Work
This section provides an overview of related work on LLM for compiler.
Automatic vectorization. The research on automatic vectorization has been a hot topic. Franchetti and Püschel(franchetti2008generating, ) use the term rewriting technique to vectorize small matrix kernels. Thomas et al.(thomas2024automatic, ) use equality saturation to automatically generate high-quality vectorizing compiler for digital signal processors, creating a new idea for automatic vectorization.
LLM for compiler. In recent years, there has been a lot of work using LLMs to optimize compilers, focusing on different aspects of code work. Armengol-Estapé et al.(armengol2024slade, ) demonstrate an innovative decompilation using a small language model to ensure correctness while improving code readability. Program repair is another study that has been highlighted. Bouzenia et al.(bouzenia2024repairagent, ) present a method of using LLMs to automatically repair programs, using iterative interaction, and shows how large models interact with tools .Iterative interaction framework(bi2024iterative, )(madaan2023self, ) is widely used in LLM applications to improve the ability of generation. In addition to optimizing code directly with LLM, another idea is to generate transform pass that can optimize code. Cummins et al.(cummins2024don, ) use the way of chain of thought to generate transform pass, thus improving the accuracy of LLM optimization. In addition, the LLM is used to predict the pass list, which is another way to avoid code correctness verification(cummins2023large, ). By adjusting the pass list, the performance of the program can be greatly changed.
ML for vectorization. Machine learning techniques are now being used to improve automatic vectorization. Kevin et al.(stock2012using, ) replace the cost model of automatic vectorization with a machine learning model to help the compiler make vectorization decisions. Ameer et al.(haj2020neurovectorizer, ) deploy a reinforcement learning method in the compiler to predict the optimal vectorization factor, which improves the performance of loop vectorization. Taneja et al.(10.1145/3696443.3708929, ) propose an innovative approach to vectorization, using LLM to generate source code that uses Intrinsic instructions, thus completing vectorization at the source level and verifying the semantic correctness of the vectorized code.
6. Conclusions
We implement a code transformation (source to souce) framework based on LLMs (large language models), which can effectively transform source code into a form that is more friendly for automatic vectorization. This framework gives full play to the ability of automatic vectorization of compiler. Furthermore, it uses the unit test and formal verification tools to ensure the correctness of generated code. This combination of the LLM and compiler reduces the responsibility of LLM, thus decreasing the probability of error. The proposed framework makes full use of the inference ability of LLMs, while preserving the accuracy of the traditional compilation vectorization. This combination will provide hints for future research. In future work, we will further ensure that the generated code is correct and extend the LLM to more compilation optimizations.